OpenStudy (df001):

Is Cot=1/tanƟ the same as Cot=CosƟ/SinƟ? My friend told me that they both work the same way when substituting in a Trig equation.

OpenStudy (anonymous):

Yes, this is the definition of cotangent. \[\cot\theta=\frac{1}{\tan\theta}\] Recall that \(\tan\theta\) is defined as follows, \[\tan\theta=\frac{\sin\theta}{\cos\theta}\] This means then that \[\cot\theta=\frac{1}{\tan\theta}=\frac{1}{\frac{\sin\theta}{\cos\theta}}=\frac{\cos\theta}{\sin\theta}\]

OpenStudy (df001):

When I substitue in a Trigonometric equation, I can use either one of the reciprocal identity or the quotient identity?

OpenStudy (anonymous):

Yep, identities are identities are identities are identities...

OpenStudy (df001):

Thanks sir

OpenStudy (anonymous):

yw